jhonyy9 (jhonyy9):

how we can prove the irrationality of number ,,pi" pi = 3,14....... my idea by reductio ad absurdum method suppose pi = a/b so what mean that b*pi = a but we know that than we multiplie the pi by indifferent integer b so never not will result an integer a your opinion please ???

jhonyy9 (jhonyy9):

@Kainui @TheSmartOne

jhonyy9 (jhonyy9):

@Directrix your opinion please about this ? thank you

jhonyy9 (jhonyy9):

@zepdrix

jhonyy9 (jhonyy9):

@satellite73 please your opinion about this ? thank you

OpenStudy (kainui):

Yeah might be possible.

OpenStudy (kainui):

 Here's a thought, since \(\pi = \frac{C}{d}\) we can at maybe try to show C is not a multiple of d. From this picture I have set up the inequality that going from left to right we can either take the straight path, the half circle or the half square. That lets us set up the inequality: \[d < \frac{C}{2} < 2d\] Which lets us say for sure: \[2 < \pi < 4\] Haha maybe not super productive. I was thinking perhaps by showing by some kind of infinite descent by bounding below and above by regular polygon paths that we'd end up somehow showing it perhaps?